# UniHetero: Could Generation Enhance Understanding for Vision-Language-Model at Large Data Scale?

### v1

@@ Abstract
UniHetero: Could Generation Enhance Understanding for Vision-Language-Model at Large Data Scale?
Fengjiao Chenthanks: Most of the work has been done before 2025.3. Please feel free to contact chenfengjiao02@meituan.com for detailed discussion or cooperation.   Minhao Jing   Weitao Lu   Yan Feng   Xiaoyu Li   Xuezhi Cao Meituan, Beijing, China.
@@ 1 Introduction
In unified vision models, by integrating understanding tasks and generation tasks, it is "expected" that generation and understanding mutually reinforce each other, demonstrating a broader range of task formats and higher task performance. The term "expected" is used because the mutual enhancement between understanding tasks and generation tasks has not yet been fully validated, especially regarding the promotion of understanding through generation.
Existing studies have validated the positive conclusion of "generation enhances understanding" in small-scale experimental settings  (Tong et al., 2024; Wang et al., 2024a). When the data scale is small, the enhancement brought by the data itself may have a greater impact compared to the task format. However, in large-scale production scenarios (with training data exceeding 100M), the effect of generation tasks on understanding tasks is often negative (Team, 2025; Wang et al., 2024b). Several works (Wu et al., 2024; Ma et al., 2025a) discuss one of the reasons for the difficulty in integrating the two tasks: visual encoding differences. Understanding tasks require high-level semantic representations, while generation tasks require low-level texture representations. The disparity between these representations makes it challenging for existing visual encodings to accommodate both simultaneously. Improvements (Ma et al., 2025a) attempt to fuse these two representations into a unified visual encoder, which shifts the difficulty from the LLM to the encoder which is challenging to solve as well.
To determine whether generation tasks can enhance understanding tasks, it is first essential to validate this in large-scale settings (>100M). In this work, we propose a concise autoregressive model with continuous representation and conduct a series of ablation experiments that verifies "generation promotes understanding" over 200M samples with better data-scale manner. Main contributions
  • Demonstrates that generation enhances understanding on over 200M samples by relying on visual semantic representations rather than raw pixel-level representations.
@@ 2.1 Unified Vision Language Models
Vision-language large models are moving toward the unification of visual understanding and visual generation tasks. Existing unified vision models can be structurally categorized into end-to-end architectures and concatenated architectures. In end-to-end architectures, both understanding and generation tasks use the same visual encoder to feed visual information into the LLM backbone. The LLM’s parameters adapt to emphasize different tasks depending on the data, thereby achieving comprehensive task fusion (Team, 2025; Wang et al., 2024b; Tong et al., 2024; Ge et al., 2025). In concatenated architectures, distinct visual encoders are used for understanding and generation tasks to feed visual information into the LLM backbone. The LLM’s parameters are explicitly partitioned by task, which prevents performance deterioration on individual tasks, but this design comes at the expense of limiting the possibility of task fusion (Wu et al., 2024; Deng et al., 2025). Additionally, there are unified models optimized specifically for generation tasks, aiming to enhance generation through understanding without preserving the performance of understanding tasks (Xiao et al., 2024; Tian et al., 2025; Zhou et al., 2024).
@@ 2.2 Understanding Enhancement Challenge
Existing work has validated that understanding enhances generation (Tian et al., 2025; Xiao et al., 2024), but there is a lack of effective verification for generation enhancing understanding, especially under large-scale data pretraining. MetaMorph (Tong et al., 2024), with a sample size of 1M and training only during the SFT stage, demonstrated that as generation data increases, the metrics for understanding tasks show an upward trend. ROSS (Wang et al., 2024a), pretrained on less than 1M samples, concluded through ablation studies that generation data promotes understanding tasks. When the data scale is small, the impact of the data itself on improvement is greater than the task format. However, in large-scale production scenarios (training data volume greater than 100M), the effect of generation tasks on understanding tasks is mostly negative (Team, 2025; Wang et al., 2024b). Prior studies (Wu et al., 2024; Ma et al., 2025a) point out that a key challenge in unifying these two tasks lies in their differing visual encodings: understanding tasks depend on high-level semantic features, whereas generation tasks require low-level texture details. This mismatch in required representations makes it difficult for current visual encoders to effectively support both types of tasks at once. UniTok (Ma et al., 2025a) attempted to fuse these two representations by enhancing representational capabilities through multi-level codebooks, validating that generation can promote understanding under discrete encoding with a sample size of 30M.
Leading visual models in understanding primarily use continuous encoding, whereas unified visual encoders mainly rely on discrete encoding. By distilling existing continuous encoders to obtain semantic representations, they can approximate but hardly surpass them. At the same time, unifying visual representations is inherently challenging, effectively shifting the task fusion problem onto the visual encoder. To further explore the improvement of understanding effects through the promotion of generation tasks based on existing visual understanding models, this work utilizes existing continuous encoders to construct a concise model structure, validating a better data-scale trend under large-scale pretraining.
@@ 3.1 Model Structure
To investigate the impact of generation tasks on understanding tasks, we propose integrating generation tasks into a visual understanding model while maintaining a concise architecture. This approach aims to preserve visual understanding capabilities while obtaining a multimodal autoregressive model. To mitigate conflicts among representations across different dimensions (in Section.2.2), we intend to fuse multimodal semantic representations inside the LLM and perform modality-specific generation through dedicated "decoders" for each modality, as shown in Figure.1.
@@
Based on the aforementioned considerations, we adopt the model architecture illustrated in Figure.2. To better preserve comprehension capabilities, unlike (Team, 2025; Wang et al., 2024b), we employ continuous encoding to represent visual information, where VAE (Li et al., 2024) represents visual generative representations (noted as pixel feature), and DINOv2 (Oquab et al., 2024) represents visual understanding representations (noted as semantic feature). At the LLM layer, only DINOv2 is used alongside text to perform multimodal auto-regression in the semantic dimension. In the vision decoder component, the image tokens output by the LLM serve as conditional inputs to the diffusion structure, which completes the details, decodes the image VAE features, and ultimately yields pixel-level images.
Consider the visual token with semantic representation $x^{s}\in\mathbb{R}^{ds}$ and pixel-level representation $x^{p}\in\mathbb{R}^{dp}$, which denotes the ground-truth token to be predicted. The LLM takes the aligned vector $e\in\mathbb{R}^{D}$ from the visual projector and generates the last hidden states $z\in\mathbb{R}^{D}$ at that same position. For semantic autoregression, we have Equation. 1, which is also used in Seed-X (Ge et al., 2025) and MetaMorph (Tong et al., 2024)).
$\begin{split}\mathcal{P}(x^{s}_{i}\mid x^{s}_{<i})&=\mathcal{P}(x^{s}_{i}\mid z_{i})\,\mathcal{P}(z_{i}\mid x^{s}_{<i})\\[4.0pt] &=\mathcal{P}(x^{s}_{i}\mid z_{i})\,\mathcal{P}(z_{i}\mid e_{<i})\,\mathcal{P}(e_{<i}\mid x^{s}_{<i})\end{split}$
@@ 3.2 Training Details
For pretraining, we employ an internal corpus of 80 million image–text aligned pairs tailored for image understanding, running for 3 epochs to yield 240 million training instances in total. Throughout training, images are randomly positioned either before or after the text, with the latter arrangement treated as image generation data. Finetuning is then performed on a small internal dataset to initially elicit the model’s capabilities and support experimental evaluation.
We employ Llama2-7B (Touvron et al., 2023) as the LLM backbone and adopt the training configurations from Chameleon (Team, 2025) to facilitate modality fusion, including techniques such as qknorm and zloss. For image representations, semantic features are encoded using DINOv2-large (Oquab et al., 2024), while pixel-level features are encoded from MAR-KL16 (Li et al., 2024). Both vision encoders are frozen during training to minimize confounding factors. The semantic features are integrated into the LLM via full-attention mechanisms, with masking probabilities sampled from a Gaussian Distribution (mean = 0.7, detailed in Section 3.3.2). As for vision decoder, we follow the similar way in MAR (Li et al., 2024) for training and inference, where the pixel-level features are decoded using flow matching.
@@ 4.1 Generation Enhances Understanding
Under the UniHetero model, we analyze whether "generation enhances understanding" through ablation of the image generation loss. The experiment settings are shown in the Table. 1. The control group exp1 uses only the text generation loss, while the experimental groups exp2 and exp3, adds two types of image generation losses on top of the control group respectivelly. In terms of implementation, exp2 does not completely remove the diffusion loss but prevents gradients from back propagating to the LLM, thereby reducing the impact on image understanding while retaining image generation capabilities.
@@
Table 1: Ablation performences of UniHetero on various image generation strategies at 240M data-scale.
@@
Furthermore, introducing diffuLoss causes a decline in understanding performance and weakens the scaling trend, as illustrated in Figure 5. A conflict between semantic representation and pixel-level representation emerges, which is consistent with findings reported in previous studies (Chen et al., 2025; Wu et al., 2024). It seems that the vision decoder module accomplishes the representation transformation through leveraging LLM rather than by itself.
Could the transformation decrease the dependence on LLMs? To explore this further, an additional experiment (denoted "+warmup") was performed by increasing the size of the diffusion structure and applying a warm-up phase to the model parameters. In practice, we adopt the pretrained mar-base-f32 model (Li et al., 2024), which is concise than Stable Diffusion (Esser et al., 2024) and shows remarkable class-condition generation performance. The experimental findings show a substantial boost in performance on understanding tasks, suggesting that image generation can potentially be realized without relying on LLM-based compensation or intricate architectures.
Figure 4: Performance of UniHetero with ploss only under large data-scale. The dashed lines indicates linear regression of data scaling law.
Figure 5: Ablation performance on image generation strategies. “+ploss” denotes UniHetero with only ploss. “+diffuloss” denotes UniHetero with both ploss and diffuloss. “+warmup” denotes UniHetero with ploss, diffuloss, and a warmed diffusion module.
@@ 4.2 Module Selection Analysis
As shown in Table. 2, we first compare mlp-cos, diffu-mse and mlp-mse on caption tasks for a quick observation. Results show that mlp-cos performs the best, which aligns with conclusions from existing studies (Tong et al., 2024; Wang et al., 2024a; Ge et al., 2025).
Next, we conduct finer experiments with end-to-end metrics (MMBench (Liu et al., 2024) and SeedBench (Li et al., 2023)) to select the module setting. Experiments indicate that fitting the input to the LLM layer yields superior results compared to fitting the output of the Vision Encoder. Specifically, the ranking observed is: ema-mlp-llm-cos > norm-3-mlp-cos > mlp-cos. It verifies the hypothesis in Equation. 2, i.e. autoregression on LLM input embedding shows less cumulative error and better modality adaptation. Compared to the commonly used fitting vision encoder (Equation. 1), it needs more parameters (in norm-3-mlp-cos setting) to learn a comparable performance.
@@ 4.3 Visualize Semantic Representation
In the previous section’s experiments, we observed that through autoregressive modeling of visual-semantic representations, generation can be verified to facilitate understanding on a large-scale dataset. Then, do visual-semantic representations actually capture? To investigate this, we now turn to a qualitative analysis.
@@ Appendix B Generation Improvement on Inference Stage
The experimental results are shown in the Figure.9. For most generated samples, quality improves after one random iteration, with issues such as distortion and blurry details being mitigated. However, there are cases where the original generated image was of good quality but degrades after regeneration, indicating that selective regeneration is still necessary. This aligns with the concept of "inference-time scaling" in the generation field (Ma et al., 2025b; Guo et al., 2025), suggesting that this direction has potential for further exploration.

### v2

@@ 1 Introduction
In unified vision models, by integrating understanding tasks and generation tasks, it is ”expected” that generation and understanding mutually reinforce each other, demonstrating a broader range of task formats and higher task performance. The term ”expected” is used because the mutual enhancement between understanding tasks and generation tasks has not yet been fully validated, especially regarding the enhancement of understanding through generation.
Existing studies have validated the positive conclusion of ”generation enhances understanding” in small-scale experimental settings  (tong2024metamorph; wang2024ross). When the data scale is small, the enhancement brought by the data itself may have a greater impact compared to the task format. However, in large-scale production scenarios (with training data exceeding 100M), the effect of generation tasks on understanding tasks is often negative (chameleonteam2025chameleon; wang2024emu3). Several works (wu2024janus; ma2025unitok) discuss one of the reasons for the difficulty in integrating the two tasks: visual encoding differences. Understanding tasks require high-level semantic representations, while generation tasks require low-level texture representations. The disparity between these representations makes it challenging for existing visual encodings to accommodate both simultaneously. Improvements (ma2025unitok) attempt to fuse these two representations into a unified visual encoder, which shifts the difficulty from the LLM to the encoder which is challenging to solve as well.
To determine whether generation tasks can enhance understanding tasks, it is first essential to validate this in large-scale settings (¿100M). In this work, we propose a concise autoregressive model with continuous representation and conduct a series of ablation experiments that verifies ”generation enhances understanding” over 200M samples with better data-scale manner. Main contributions
  • Demonstrates that generation enhances understanding on over 200M samples by relying on visual semantic representations rather than pixel-level representations.
@@ 2.1 Unified Vision Language Models
Vision-language large models are moving toward the unification of visual understanding and visual generation tasks. Existing unified vision models can be structurally categorized into end-to-end architectures and concatenated architectures. In end-to-end architectures, both understanding and generation tasks use the same visual encoder to feed visual information into the LLM backbone. The LLM’s parameters adapt to emphasize different tasks depending on the data, thereby achieving comprehensive task fusion (chameleonteam2025chameleon; wang2024emu3; tong2024metamorph; ge2025seedx). In concatenated architectures, distinct visual encoders are used for understanding and generation tasks to feed visual information into the LLM backbone. The LLM’s parameters are explicitly partitioned by task, which prevents performance deterioration on individual tasks, but this design comes at the expense of limiting the possibility of task fusion (wu2024janus; deng2025bagel). Additionally, there are unified models optimized specifically for generation tasks, aiming to enhance generation through understanding without preserving the performance of understanding tasks (xiao2024omnigen; tian2025unigen; zhou2024transfusion).
@@ 2.2 Understanding Enhancement Challenge
Existing work has validated that understanding enhances generation (tian2025unigen; xiao2024omnigen), but there is a lack of effective verification for generation enhancing understanding, especially under large-scale data pretraining. MetaMorph (tong2024metamorph), with a sample size of 1M and training only during the SFT stage, demonstrated that as generation data increases, the metrics for understanding tasks show an upward trend. ROSS (wang2024ross), pretrained on less than 1M samples, concluded through ablation studies that generation data enhances understanding tasks. When the data scale is small, the impact of the data itself on improvement is greater than the task format. However, in large-scale production scenarios (training data volume greater than 100M), the effect of generation tasks on understanding tasks is mostly negative (chameleonteam2025chameleon; wang2024emu3). Prior studies (wu2024janus; ma2025unitok) point out that a key challenge in unifying these two tasks lies in their differing visual encodings: understanding tasks depend on high-level semantic features, whereas generation tasks require low-level texture details. This mismatch in required representations makes it difficult for current visual encoders to effectively support both types of tasks at once. UniTok (ma2025unitok) attempted to fuse these two representations by enhancing representational capabilities through multi-level codebooks, validating that generation can enhance understanding under discrete encoding with a sample size of 30M.
Leading visual models in understanding primarily use continuous encoding, whereas unified visual encoders mainly rely on discrete encoding. By distilling existing continuous encoders to obtain semantic representations, they can approximate but hardly surpass them. At the same time, unifying visual representations is inherently challenging, effectively shifting the task fusion problem onto the visual encoder. To further explore the improvement of understanding effects through the enhancement of generation tasks based on existing visual understanding models, this work utilizes existing continuous encoders to construct a concise model structure, validating a better data-scale trend under large-scale pretraining.
@@ 3.1 Model Structure
To investigate the impact of generation tasks on understanding tasks, we propose integrating generation tasks into a visual understanding model while maintaining a concise architecture. This approach aims to preserve visual understanding capabilities while obtaining a multimodal autoregressive model. To mitigate conflicts among representations across different dimensions (in Section.2.2), we intend to fuse multimodal semantic representations inside the LLM and perform modality-specific generation through dedicated ”decoders” for each modality, as shown in Figure.2.
@@
Based on the aforementioned considerations, we adopt the model architecture illustrated in Figure.1. To better preserve comprehension capabilities, unlike (chameleonteam2025chameleon; wang2024emu3), we employ continuous encoding to represent visual information, where VAE (li2024mar) represents visual generative representations (noted as pixel feature), and DINOv2 (oquab2024dinov2) represents visual understanding representations (noted as semantic feature). At the LLM layer, only DINOv2 is used alongside text to perform multimodal auto-regression in the semantic dimension. In the vision decoder component, the image tokens output by the LLM serve as conditional inputs to the diffusion structure, which completes the details, decodes the image VAE features, and ultimately yields pixel-level images.
Consider the visual token with semantic representation $x^{s}\in\mathbb{R}^{ds}$ and pixel-level representation $x^{p}\in\mathbb{R}^{dp}$, which denotes the ground-truth token to be predicted. The LLM takes the aligned vector $e\in\mathbb{R}^{D}$ from the visual projector and generates the last hidden states $z\in\mathbb{R}^{D}$ at that same position. For semantic autoregression, we have Equation. 1, which is also used in Seed-X (ge2025seedx) and MetaMorph (tong2024metamorph)).
$\begin{split}\mathcal{P}(x^{s}_{i}\mid x^{s}_{<i})&=\mathcal{P}(x^{s}_{i}\mid z_{i})\,\mathcal{P}(z_{i}\mid x^{s}_{<i})\\[4.0pt] &=\mathcal{P}(x^{s}_{i}\mid z_{i})\,\mathcal{P}(z_{i}\mid e_{<i})\,\mathcal{P}(e_{<i}\mid x^{s}_{<i})\end{split}$ (1)
@@ 3.2 Training Details
For pretraining, we employ an internal corpus of 80 million image–text aligned pairs tailored for image understanding, running for 3 epochs to yield 240 million training instances in total. Training batch size is 64 and the iteration amount is 37.5K. Throughout training, images are randomly positioned either before or after the text, with the latter arrangement treated as image generation data. Finetuning is then performed on a small internal dataset to initially elicit the model’s capabilities and support experimental evaluation.
We employ Llama2-7B (touvron2023llama2) as the LLM backbone and adopt the training configurations from Chameleon (chameleonteam2025chameleon) to facilitate modality fusion, including techniques such as qknorm and zloss. For image representations, semantic features are encoded using DINOv2-large (oquab2024dinov2), while pixel-level features are encoded from MAR-KL16 (li2024mar). Both vision encoders are frozen during training to minimize confounding factors. The semantic features are integrated into the LLM via full-attention mechanisms, with masking probabilities sampled from a Gaussian Distribution (mean = 0.7, detailed in Section 3.3.2). As for vision decoder, we follow the similar way in MAR (li2024mar) for training and inference, where the pixel-level features are decoded using flow matching.
@@ 4.1 Generation Enhances Understanding
Under the UniHetero model, we analyze whether ”generation enhances understanding” through ablation of the image generation loss. The experiment settings are shown in the Table. 1. The control group exp1 uses only the text generation loss, while the experimental groups exp2 and exp3, adds two types of image generation losses on top of the control group respectivelly. In terms of implementation, exp2 does not completely remove the diffusion loss but prevents gradients from back propagating to the LLM, thereby reducing the impact on image understanding while retaining image generation capabilities.
@@
Table 1: Ablation performences of UniHetero on various image generation strategies at more than 200M data-scale.
@@
Furthermore, introducing diffuLoss causes a decline in understanding performance and weakens the scaling trend, as illustrated in Figure 5. A conflict between semantic representation and pixel-level representation emerges, which is consistent with findings reported in previous studies (chen2025blip3o; wu2024janus). It seems that the vision decoder module accomplishes the representation transformation through leveraging LLM rather than by itself.
Could the transformation decrease the dependence on LLMs? To explore this further, an additional experiment (denoted ”+warmup”) was performed by increasing the size of the diffusion structure and applying a warm-up phase to the model parameters. In practice, we adopt the pretrained mar-base-f32 model (li2024mar), which is concise than Stable Diffusion (esser2024stablediffu) and shows remarkable class-condition generation performance. The experimental findings show a substantial boost in performance on understanding tasks, suggesting that image generation can potentially be realized without relying on LLM-based compensation or intricate architectures.
Figure 4: Performance of UniHetero with ploss using over 200M data samples. The x-axis represents the training steps. The dashed lines show the linear regression fit for the data scaling law.
Figure 5: Ablation results for different vision generation configurations. “+ploss” indicates UniHetero using only ploss. “+diffuloss” refers to UniHetero using both ploss and diffuloss. “+warmup” corresponds to UniHetero employing ploss, diffuloss, and a warmed-up diffusion module.
@@ 4.2 Module Selection Analysis
As shown in Table. 2, we first compare mlp-cos, diffu-mse and mlp-mse on caption tasks for a quick observation. Results show that mlp-cos performs the best, which aligns with conclusions from existing studies (tong2024metamorph; wang2024ross; ge2025seedx).
Next, we conduct finer experiments with end-to-end metrics (MMBench (liu2024mmbench) and SeedBench (li2023seedbench)) to select the module setting. Experiments indicate that fitting the input to the LLM layer yields superior results compared to fitting the output of the Vision Encoder. Specifically, the ranking observed is: ema-mlp-llm-cos ¿ norm-3-mlp-cos ¿ mlp-cos. This confirms the hypothesis in Equation. 2, namely that performing autoregression on the LLM input embeddings results in lower cumulative error and improved modality adaptation. To achieve a similar level of performance, the commonly adopted fitting vision encoder (Equation. 1) requires a larger number of parameters under the norm-3-mlp-cos configuration.
@@ 4.3 Visualize Semantic Representation
In the previous section’s experiments, we observed that through autoregressive modeling of visual-semantic representations, generation can be verified to improve understanding on a large-scale dataset. Then, do visual-semantic representations actually capture? To investigate this, we now turn to a qualitative analysis.
@@ 5 Conclusion
Acknowledgement
We hereby express our appreciation to the LongCat Team EVA Committee for their valuable assistance, guidance, and suggestions throughout the course of this work.
@@ Appendix B Generation Improvement on Inference Stage
The experimental results are shown in the Figure.9. For most generated samples, quality improves after one random iteration, with issues such as distortion and blurry details being mitigated. However, there are cases where the original generated image was of good quality but degrades after regeneration, indicating that selective regeneration is still necessary. This aligns with the concept of ”inference-time scaling” in the generation field [ma2025inferencetimescaling, guo2025generateimagescotlets], suggesting that this direction has potential for further exploration.